Andy’s Accessories produces backpacks. The costs and prices for the backpacks follow (Assume the same unit costs in all years):

Selling price                                          $23.00 per backpack
Variable costs:
Production $11.00 per backpack
Selling $2.00 per backpack
Fixed Costs:
Production                                           $900,000 per year
Selling and administrative $540,000 per year

Andy’s Accessories produced 250,000 backpacks for the year and sold 200,000. There was no beginning inventory, and costs throughout the year were stable. How much is the cost of ending inventory under full costing?

Group of answer choices

$550,000

$938,000

$650,000

$730,000
Answer #1

Cost per unit under full costing

= Variable production cost per unit + Fixed production cost per unit

= 11 + (900,000/250,000) = 11+3.60 = 14.60 per unit

Cost of ending inventory

= (250,000-200,000)*14.60

= 50,000*14.60

= $730,000